News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Tuesday 3 August 1999 7.00am EST

The Chandra mission continues to proceed nominally.

The following is an overview of the events that occurred over the last 24 hours.

Continued ACIS calibration using internal sources.

Develop and verify all command loads and software patches to implement IPS Burn # 4. Continued review of data from IPS Burn #3,and ground test data from firings conducted during LAE development. The objective of this activity being to more precisely determine the cause of engine #3 under performance during IPS burn #3.

Earth Scan 17 is in process.

Events scheduled for the next 24 hours
  • ACIS Long Form Functional Test
  • Continuation of HRC Checkout Activities
  • Final review with Chandra program, NASA and TRW management of decision to use LAE's 2 & 4 for IPS #4.
  • Dry run IPS Burn # 4
  • Earth Scan 18

Return to Status Reports